Please help me : Find the standard form of the equation of the line that passes through point (1,-4) and is parallel to the line 4x-y=2
Answer by stanbon(75887)

Please help me : Find the standard form of the equation of the line that passes through point (1,-4) and is parallel to the line 4x-y=2
---
Solve for "y" to find the slope of the given line:
y = 4x-2
slope = 4
----------------
Form:
y = mx + b
-4 = 4*1+b
b = -8
------------
Equation:
y = 4x-8
----------------------
Cheers,
Stan H.
=============